OpenStudy (anonymous):

A plant 3 in. tall grows an average of 0.5 in. each month. Which equation models the flower’s height h after x months?

OpenStudy (anonymous):

I think the equation is f(x)=3(0.5)^x

OpenStudy (♪chibiterasu):

Plant = p Inches = n Month = x Height = h h = p(ix) h = 3(0.5 • x) @greenlegodude57 Equation, not function.

OpenStudy (anonymous):

Oh, I see.